OpenStudy (anonymous):

Sandra is riding the Ferris wheel, and her height can be modeled by the equation H(t) = 25 cospi over 14t + 31, where H represents the height of the person above the ground in feet at t seconds. Part 1: How far above the ground is Sandra before the ride begins? Part 2: How long does the Ferris wheel take to make one complete revolution? Part 3: Assuming Sandra begins the ride at the top, how far from the ground is the edge of the Ferris wheel when Sandra's height above the ground reaches a minimum? please help asap

OpenStudy (anonymous):

part one put \(t=0\) since that is when you start

OpenStudy (anonymous):

i can't really read it is it\[\large 25\cos(\frac{\pi}{14}t)+31\]?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

weird because that means she gets on a the highest point

OpenStudy (anonymous):

i dont get this at all

OpenStudy (anonymous):

put \(t=0\) and \(\cos(0)=1\) so you get \[25\times 1+31=56\]

OpenStudy (anonymous):

" H represents the height of the person above the ground in feet at t seconds." that meas for example at time \(t=0\) which is when you start counting, replace \(t\) by \(0\) like i just did above

OpenStudy (anonymous):

then after say \(7\) seconds the height would be \[25\cos(\frac{7}{14}\pi)+31=25\cos(\frac{\pi}{2})+31=0+31=31\] since \(\cos(\frac{\pi}{2})=0\)

OpenStudy (anonymous):

Is this still for the first part?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

part one just asks for \(t=0\)

OpenStudy (anonymous):

which is \(56\) since \(25+31=56\)

OpenStudy (anonymous):

okay but for part 2 how do you find the complete revolution

OpenStudy (anonymous):

that is the same as the period of the function

OpenStudy (anonymous):

the period is always \[\frac{2\pi}{b}\] and in your case \(b=\frac{\pi}{14}\) how is your arithmetic?

OpenStudy (anonymous):

so 28

OpenStudy (anonymous):

that was a serious question, not a snark many cannot do this yes 28

OpenStudy (anonymous):

and for part 3 if she starts at the top and ends at the bottom that would be 180°????????

OpenStudy (anonymous):

well not since you are working in radians, no

OpenStudy (anonymous):

what you really mean is when cosine is minus one (its minimum) so don't think too hard, just replace cosine by minus one

OpenStudy (anonymous):

i understand you meant put \(t=180\) which won't work here since the measure is in radians the idea is right, but really all you are asked is the minimum value of \[25\cos(whatver)+31\] which is \[-25+31\] since \(-1\) is the minimum value of cosine

OpenStudy (anonymous):

H(t)=25(-1)(╥/14)t+31

OpenStudy (anonymous):

and just solve

OpenStudy (anonymous):

no no just replace cosine by minus one forget the input

OpenStudy (anonymous):

you are thinking way too hard the maximum value of cosine is 1 so the max of \[25\cos(x)+31\] is \[25+31=56\] like in question one the minumum of cosine is \(-1\) so the minimum of \[25\cos(x)+31\] is \[-25+31=6\]

OpenStudy (anonymous):

H(t)=6(╥/14)t ?

OpenStudy (anonymous):

what about ╥/14

OpenStudy (anonymous):

you are not asked "When" is it the minimum only "what is the minimum"

OpenStudy (anonymous):

╥=pi

OpenStudy (anonymous):

yeah i got that if you were asked "when is it the minimum" then you would have to worry about that

OpenStudy (anonymous):

oh okay can you help me with one more?

OpenStudy (anonymous):

but you are only asked "what is the minimum" and that you can do in your head it is \(-25+31=6\)

OpenStudy (anonymous):

sure if i can

OpenStudy (anonymous):

thanks Part 1: Using complete sentences, compare the key features and graphs of sine and cosine. What are their similarities and differences? Part 2: Using these similarities and differences, how would you transform f(x) = 3 sin(4x - π) + 4 into a cosine function in the form f(x) = a cos(bx - c) + d?

OpenStudy (anonymous):

oh damn the dreaded "complete sentences"

OpenStudy (anonymous):

haha yea

OpenStudy (anonymous):

if it was me, i would say the graphs are identical they both have max an min 1 and -1 they are both periodic with period \(2\pi\) the only difference is that say compared to the graph of sine, the graph of cosine is shifted \(\frac{\pi}{2}\) units to the left

OpenStudy (anonymous):

okay i get that but what about the second part?

OpenStudy (anonymous):

ok since \[\sin(x+\frac{\pi}{2})=\cos(x)\] perhaps we could translate that way

OpenStudy (anonymous):

the 3 out front will have to be the same, also the \(+4\) and the end will have to be the same so it will look like \[3\cos(something)+4\] all we need is the "something"

OpenStudy (anonymous):

so how do we find it

OpenStudy (anonymous):

well i am sure there is a slicker way to do it, but i will do it the donkey way lets see what makes \(4x-\pi=0\)

OpenStudy (anonymous):

we get \(x=\frac{\pi}{4}\) pretty much right away

OpenStudy (anonymous):

so when \(x=\frac{\pi}{4}\) we get 0 and the sine of 0 is 0 now the cosine of \(-\frac{\pi}{2}=0\) so we want \[bx-c=-\frac{\pi}{4}\] when ' \[x=\frac{\pi}{4}\]

OpenStudy (anonymous):

\[b\frac{\pi}{4}-c=-\frac{\pi}{4}\] we have lots of choices, since this is one equation and two unknowns so lets make \(b=4\) and get \[\pi-c=-\frac{\pi}{4}\] making \(c=\frac{5\pi}{4}\) if my Lagrange is right

OpenStudy (anonymous):

i am sure there is a slicker way to do it, but i think we end up with \[3\cos(4x-\frac{5\pi}{4})+4\]

OpenStudy (anonymous):

Thank you!!!!!!!!!!!!!!!

OpenStudy (anonymous):

hold the phone i am wrong!!

OpenStudy (anonymous):

i shifted incorrectly some how give me a minute to figure out my mistake

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

somewere i made an algebra mistake it should be \[3\cos(4x+\frac{\pi}{2})+4\] or

OpenStudy (anonymous):

jeez i screwed that up \[bx-c=-\frac{\pi}{4}\] was a mistake, it should have been \[bx-c=-\frac{\pi}{2}\]

OpenStudy (anonymous):

should i replace all the pi/4 to pi/2

OpenStudy (anonymous):

when \(x=\frac{\pi}{4}\) and we solve \[\pi-c=-\frac{\pi}{2}\] and get \(c=\frac{3\pi}{2}\) making it \[3\cos(4x-\frac{3\pi}{2}+4\]

OpenStudy (anonymous):

So we find what makes 4x-╥=0, we get x=╥/4, so when x=╥/4 we get 0 and the sine of 0 is 0. Now the cosine of -╥/2=0 so we want bx-c=-╥/2 when x=╥/4. b*╥/4-c=-╥/4, lets make b = 4, ╥-c=-╥/4, making c=5╥/4, so we end up with 3cos(4x+3╥/2)+4.

OpenStudy (anonymous):

the part where i solved \[4x-\pi=0\] and got \[x=\frac{\pi}{4}\] is still correct

OpenStudy (anonymous):

is that okay?

OpenStudy (anonymous):

╥=pi

OpenStudy (anonymous):

So we find what makes 4x-╥=0, we get x=╥/4, so when x=╥/4 we get 0 and the sine of 0 is 0. Now the cosine of -╥/2=0 so we want bx-c=-╥/2 when x=╥/4. b*╥/4-c=-╥/2, lets make b = 4, ╥-c=-╥/2, making c=3╥/2, so we end up with 3cos(4x-3╥/2)+4.

OpenStudy (anonymous):

there is the corrected version

OpenStudy (anonymous):

i changed my mistaken \(\frac{\pi}{4} s\) to the correct \(\frac{\pi}{2}\)

OpenStudy (anonymous):

i am sure there is a slicker way to do it ,but i can't think of it at the moment

OpenStudy (anonymous):

its okay thank u very much

OpenStudy (anonymous):

yw